# The revised boomerang connectivity tables and their connection to the difference distribution table

### v2

@@ 2. $(2)$
    Let $Z=X+X^{2^{s}}$. Then $F_{3}(x)+F_{3}(X+1)=c$ holds if and only if $(X,Z)\in\mathbb{F}_{2^{n}}\times\mathbb{F}_{2^{n}}$ satisfies the below equations:
    (7.3) $Z^{2}+(t+1)Z+c^{2^{s}}+c^{2^{2s}}=0$
    (7.4) $X^{2}+X+Z^{2^{s}+1}+Z^{2^{s}}+z^{2}+c+1=0.$
@@
[20, Theorem 2] Given elements $u\in\mathbb{F}_{2^{2s}}\setminus\{\mathbb{F}_{2^{s}}\}$ and $v\in\mathbb{F}_{2^{n}}\setminus\{\mathbb{F}_{2^{2s}}\}$ such that
$u+u^{2^{s}}=1,v+v^{2^{2s}}=1,$
which always exist, then $X\in W$ if and only if
$X=(tu+\beta)v+\tau,$
where $t\in\mathbb{F}_{2^{s}}\setminus\{1\},\tau\in\mathbb{F}_{2^{2s}}$, and $\beta\in\mathbb{F}_{2^{s}}$ satisfies
$\beta=(t+1)^{-1}(\alpha^{2}+\alpha)+(t+1)(u^{2}+u)+1,$
for some $\alpha\in\mathbb{F}_{2^{s}}$.
@@
Let $F_{3}(X)=X^{2^{2s}+2^{s}+1}$, where $n=4s$. Then for $a,b,c,d\in\mathbb{F}_{2^{n}}^{*}$, we have
$\operatorname{EBCT}_{F_{3}}(a,b,c,d)=\begin{cases}\operatorname{DDT}_{F_{3}}(c,d)&\mbox{ if }a=c,b=d,\\ &\mbox{ or }\operatorname{DDT}_{F_{3}}(c,d)=4,a=c((t+1)u+\alpha),\mbox{ and }\\ &\qquad b=d(\alpha+(t+1)u^{2^{s}})+F_{3}(c)\gamma,\\ &\mbox{ or }\operatorname{DDT}_{F_{3}}(c,d)=4,a=c((t+1)u+\alpha+1),\mbox{ and }\\ &\qquad b=d(\alpha+(t+1)u^{2^{s}}+1)+F_{3}(c)\gamma,\\ 0&\mbox{otherwise,}\end{cases}$
where $t={\rm Tr}_{s}^{n}\left(\frac{d}{F_{3}(c)}\right)$, and $F_{3}(tu+\beta)v+\tau)+F_{3}(tu+\beta)v+\tau+1)=\frac{d}{F_{3}(c)},$
$\operatorname{LBCT}_{F_{3}}(a,b,c)=\begin{cases}\operatorname{DDT}_{F_{3}}(b,c)&\mbox{if}~{}\operatorname{DDT}_{F_{3}}(b,c)=2\mbox{ and }a=b,\\ &\mbox{ or }\operatorname{DDT}_{F_{3}}(b,c)=4,\mbox{ and}\\ &\qquad a\in\{b,b((t+1)u+\alpha),b((t+1)u+\alpha+1)\},\\ 0&\mbox{otherwise,}\end{cases}$
where $t={\rm Tr}_{s}^{n}\left(\frac{c}{F_{3}(b)}\right)$, and $F_{3}(tu+\beta)v+\tau)+F_{3}(tu+\beta)v+\tau+1)=\frac{c}{F_{3}(b)},$ and,
$\operatorname{UBCT}_{F_{3}}(a,b,c)=\begin{cases}\operatorname{DDT}_{F_{3}}(a,b)&\mbox{if}~{}\operatorname{DDT}_{F_{3}}(a,b)=2\mbox{ and }c=b,\\ &\mbox{ or }\operatorname{DDT}_{F_{3}}(a,b)=4\mbox{ and }\\ &\qquad c\in\{b,b(\alpha+(t+1)u^{2^{s}})+F_{3}(a)\gamma,\mbox{ and}\\ &\qquad b(\alpha+(t+1)(1+u)+1)+F_{3}(a)\gamma\},\\ 0&\mbox{otherwise,}\end{cases}$
where $t={\rm Tr}_{s}^{n}\left(\frac{b}{F_{3}(a)}\right)$, and $F_{3}(tu+\beta)v+\tau)+F_{3}(tu+\beta)v+\tau+1)=\frac{b}{F_{3}(a)}.$
In all the tables computed, $\gamma=((t+1)^{-1}(\alpha^{4}+\alpha^{2})+(t+1)^{3}(u^{4}+u^{2})+\alpha^{2}t+\alpha t+t(t+1)(u+u^{2}))$, $\tau\in\mathbb{F}_{2^{2s}}$, $u\in\mathbb{F}_{2^{2s}}\setminus\{\mathbb{F}_{2^{s}}\}$, $v\in\mathbb{F}_{2^{n}}\setminus\{\mathbb{F}_{2^{2s}}\}$ satisfying $u+u^{2^{s}}=1$, $v+v^{2^{2s}}=1$ ,and $\beta\in\mathbb{F}_{2^{s}}$ satisfying $\beta=(t+1)^{-1}(\alpha^{2}+\alpha)+(t+1)(u^{2}+u)+1$, for some $\alpha\in\mathbb{F}_{2^{s}}$.
@@
$\begin{cases}\operatorname{DDT}_{F_{3}}(c,d)&\mbox{ if }a=c,b=d,\mbox{ or}\\ &\quad\operatorname{DDT}_{F_{3}}(c,d)=4,a=z_{1}+z_{2},b=F_{3}(z_{1})+F_{3}(z_{2}),\mbox{ or}\\ &\quad\operatorname{DDT}_{F_{3}}(c,d)=4,a=z_{1}+z_{2}+c,b=F_{3}(z_{1})+F_{3}(z_{2}+c)\\ 0&\mbox{otherwise,}\end{cases}$
@@
Case 2. Let $\frac{z_{2}}{c}=\frac{z_{1}}{c}+(t+1)u+\alpha+1$, then by same argument as in Case 1, we have $a=c((t+1)u+\alpha+1)$ and $b=d(\alpha+(t+1)u^{2^{s}}+1)+F_{3}(c)((t+1)^{-1}(\alpha^{4}+\alpha^{2})+(t+1)^{3}(u^{4}+u^{2})+\alpha^{2}t+\alpha t+t(t+1)(u+u^{2})).$
@@
We again use (also its notations) Corollary 7.2 to compute the LBCT and UBCT entries for the Bracken-Leander function. As we have computed $\operatorname{EBCT}_{F_{3}}(a,b,c,d)=2$, when $a=c,b=d$ and $\operatorname{DDT}_{F_{3}}(c,d)=2$, then $\operatorname{LBCT}_{F_{3}}(a,c,d)=2$, for $a=c$ when $\operatorname{DDT}_{F_{3}}(c,d)=2$ and, $\operatorname{UBCT}_{F_{3}}(c,d,b)=2$, for $b=d$ when $\operatorname{DDT}_{F_{3}}(c,d)=2$.
@@
Via our Corollary 7.1, we can also give a very short proof for the EBCT, LBCT and UBCT entries of the inverse function for $n$ even, which is the main result in [12, 16] (for $n$ odd, see the previous section).
@@
Let $F_{4}(X)=X^{2^{n}-2}$ where $n$ is even. Then for $a,b,c,d\in\mathbb{F}_{2^{n}}^{*}$, we have
$\operatorname{EBCT}_{F_{4}}(a,b,c,d)$ =$=\begin{cases}4&\mbox{ if }~{}a=c=\frac{1}{b}=\frac{1}{d},\\ &\mbox{ or }b=\frac{1}{a},d=\frac{1}{c},\mbox{ and }(ad)^{2}+ad+1=0,\\ 2&\mbox{if}~{}a=c,b=d,d\neq\frac{1}{c}\mbox{ and }{\rm Tr}\left(\frac{1}{cd}\right)=0,\\ 0&\mbox{otherwise,}\end{cases}$
$\operatorname{LBCT}_{F_{4}}(a,b,c)$ =$=\begin{cases}4&\mbox{ if }~{}a=b=\frac{1}{c},\\ &\mbox{ or }b=\frac{1}{c},(ac)^{2}+ac+1=0,\\ 2&\mbox{if}~{}a=b\neq\frac{1}{c},{\rm Tr}\left(\frac{1}{bc}\right)=0,\\ 0&\mbox{otherwise,}\end{cases}$
$\operatorname{UBCT}_{F_{4}}(a,b,c)$ =$=\begin{cases}4&\mbox{ if }~{}b=c=\frac{1}{a},\\ &\mbox{ or }b=\frac{1}{a},(ac)^{2}+ac+1=0,\\ 2&\mbox{if}~{}c=b\neq\frac{1}{a},{\rm Tr}\left(\frac{1}{ab}\right)=0,\\ 0&\mbox{otherwise.}{}\end{cases}$
@@
If $a=c,b=d$, $d\neq\frac{1}{c}$ and ${\rm Tr}\left(\frac{1}{cd}\right)=0$, then $\operatorname{EBCT}_{F_{4}}(a,b,c,d)=\operatorname{DDT}_{F_{4}}(c,d)=2$.
If $d=\frac{1}{c}$, and without loss of generality, we let $z_{1}=0,z_{2}=\omega c$. If $a=c,b=d$, then $\operatorname{EBCT}_{F_{4}}(a,b,c,d)=\operatorname{DDT}_{F_{4}}(c,d)=4$. If $a=z_{1}+z_{2}=\omega c$ and $b=F_{4}(z_{1})+F_{4}(z_{2})=\frac{1}{\omega c}$, then $\operatorname{EBCT}_{F_{4}}(a,b,c,d)=\operatorname{DDT}_{F_{4}}(c,d)=4$. If $a=z_{1}+z_{2}+c=\omega^{2}c$ and $b=F_{4}(z_{1})+F_{4}(z_{2}+c)=\frac{1}{\omega^{2}c}$, then $\operatorname{EBCT}_{F_{4}}(a,b,c,d)=\operatorname{DDT}_{F_{4}}(c,d)=4$. Note that these last two cases can be summarized as $b=\frac{1}{a}$, $d=\frac{1}{c}$ and $(ad)^{2}+ad+1=0$. In all other cases, $\operatorname{EBCT}_{F_{4}}(a,b,c,d)=0$.
@@
Clearly, when $d=\frac{1}{c}$, we can write $A=\{(c,c,d),(\omega c,c,d),(\omega^{2}c,c,d)\bigm{|}\operatorname{DDT}_{F_{4}}(c,d)=4\}$, as, in the case $\operatorname{DDT}_{F_{4}}(c,d)=4$, $(c,c,d)\in A$, for $b=d$, $(\omega c,c,d)\in A$, for $b=\frac{1}{\omega c}$ and $(\omega^{2}c,c,d)\in A$, for $b=\frac{1}{\omega^{2}c}$. Thus, from Corollary 7.2, we have $\operatorname{LBCT}_{F_{4}}(a,c,d)=4$, for $a\in\{c,\omega c,\omega^{2}c\}$ and $d=\frac{1}{c}$. This is equivalent to saying $\operatorname{LBCT}_{F_{4}}(a,c,d)=4$ when $d=\frac{1}{c}$ and $(ad)^{2}+ad=1=0$.
In a similar way, one can write $B=\{(c,d,d),(c,d,\frac{1}{\omega c}),(c,d,b=\frac{1}{\omega c})\bigm{|}\operatorname{DDT}_{F_{4}}(c,d)=4\}$, as, in the case $\operatorname{DDT}_{F_{4}}(c,d)=4$, $(c,d,d)\in B$, for $a=c$, $(c,d,\frac{1}{\omega c})\in B$, for $a=\omega c$ and $(c,d,\frac{1}{\omega c})\in B$, for $a=\omega^{2}c$.
This completes the proof. ∎

### v3

@@ 2. $(2)$
    Let $Z=X+X^{2^{s}}$. Then $F_{3}(x)+F_{3}(X+1)=c$ holds if and only if $Z^{2}+(t+1)Z+c^{2^{s}}+c^{2^{2s}}=0$ and $X^{2}+X+Z^{2^{s}+1}+Z^{2^{s}}+Z^{2}+c+1=0$.
@@
[21, Theorem 2] Given elements $u\in\mathbb{F}_{2^{2s}}\setminus\{\mathbb{F}_{2^{s}}\}$ and $v\in\mathbb{F}_{2^{n}}\setminus\{\mathbb{F}_{2^{2s}}\}$ such that $u+u^{2^{s}}=1,v+v^{2^{2s}}=1,$ which always exist, then $X\in W$ if and only if $X=(tu+\beta)v+\tau,$ where $t\in\mathbb{F}_{2^{s}}\setminus\{1\},\tau\in\mathbb{F}_{2^{2s}}$, and $\beta\in\mathbb{F}_{2^{s}}$ satisfies $\beta=(t+1)^{-1}(\alpha^{2}+\alpha)+(t+1)(u^{2}+u)+1,$ for some $\alpha\in\mathbb{F}_{2^{s}}$.
@@
Let $F_{3}(X)=X^{2^{2s}+2^{s}+1}$ on $\mathbb{F}_{2^{n}}$, where $n=4s$. Then for $a,b,c,d\in\mathbb{F}_{2^{n}}$, we have
$\operatorname{EBCT}_{F_{3}}(a,b,c,d)=\begin{cases}2^{n}&\mbox{if }a=b=c=d=0,\\ \operatorname{DDT}_{F_{3}}(a,b)&\mbox{if }c=d=0\mbox{ and }a\neq 0,\\ \operatorname{DDT}_{F_{3}}(c,d)&\mbox{if}~{}ac\neq 0,a=c\mbox{ and }b=d;\mbox{or }a=0,b=0\mbox{ and }c\neq 0;\\ &\mbox{ or }\operatorname{DDT}_{F_{3}}(c,d)=4,a=c((t+1)u+\alpha)\mbox{ and }\\ &\qquad b=d(\alpha+(t+1)u^{2^{s}})+F_{3}(c)\gamma;\\ &\mbox{ or }\operatorname{DDT}_{F_{3}}(c,d)=4,a=c((t+1)u+\alpha+1)\mbox{ and }\\ &\qquad b=d(\alpha+(t+1)u^{2^{s}}+1)+F_{3}(c)\gamma,\\ 0&\mbox{otherwise,}\end{cases}$
where $t={\rm Tr}_{s}^{n}\left(\frac{d}{F_{3}(c)}\right)$, and $F_{3}((tu+\beta)v+\tau)+F_{3}((tu+\beta)v+\tau+1)=\frac{d}{F_{3}(c)},$
$\operatorname{LBCT}_{F_{3}}(a,b,c)=\begin{cases}2^{n}&\mbox{if }b=c=0,\\ \operatorname{DDT}_{F_{3}}(b,c)&\mbox{if }a=0\mbox{ and }b\neq 0;\mbox{ or }\\ &~{}\operatorname{DDT}_{F_{3}}(b,c)=2,ab\neq 0\mbox{ and }a=b;\mbox{ or }\\ &~{}\operatorname{DDT}_{F_{3}}(b,c)=4,ab\neq 0\mbox{ and}\\ &\qquad a\in\{b,b((t+1)u+\alpha),b((t+1)u+\alpha+1)\},\\ 0&\mbox{otherwise,}\end{cases}$
where $t={\rm Tr}_{s}^{n}\left(\frac{c}{F_{3}(b)}\right)$, and $F_{3}((tu+\beta)v+\tau)+F_{3}((tu+\beta)v+\tau+1)=\frac{c}{F_{3}(b)},$ and,
$\operatorname{UBCT}_{F_{3}}(a,b,c)=\begin{cases}|F_{3}^{-1}(c+{\rm Im}(F_{3}))|&\mbox{if}~{}a=b=0,\\ \operatorname{DDT}_{F_{3}}(a,b)&\mbox{if}~{}\operatorname{DDT}_{F_{3}}(a,b)=2,a\neq 0\mbox{ and }c=b;\mbox{ or }\\ &~{}\operatorname{DDT}_{F_{3}}(a,b)=4,a\neq 0,\\ &~{}c\in\{b,b(\alpha+(t+1)u^{2^{s}})+F_{3}(a)\gamma,\mbox{ and}\\ &\qquad b(\alpha+(t+1)(1+u)+1)+F_{3}(a)\gamma\},\\ 0&\mbox{otherwise,}\end{cases}$
where $t={\rm Tr}_{s}^{n}\left(\frac{b}{F_{3}(a)}\right)$, $F_{3}((tu+\beta)v+\tau)+F_{3}((tu+\beta)v+\tau+1)=\frac{b}{F_{3}(a)},$ ${\rm Im}(F_{3})$ is the image of $F_{3}$ and $F_{3}^{-1}(\cdot)$ denotes the preimage of the argument. In particular, if $F_{3}$ is a permutation, or equivalently if $s$ is odd, then $|F_{3}^{-1}(c+{\rm Im}(F_{3}))|=2^{n}$.
In all the computed tables, $\gamma=((t+1)^{-1}(\alpha^{4}+\alpha^{2})+(t+1)^{3}(u^{4}+u^{2})+\alpha^{2}t+\alpha t+t(t+1)(u+u^{2}))$, $\tau\in\mathbb{F}_{2^{2s}}$, $u\in\mathbb{F}_{2^{2s}}\setminus\{\mathbb{F}_{2^{s}}\}$, $v\in\mathbb{F}_{2^{n}}\setminus\{\mathbb{F}_{2^{2s}}\}$ satisfying $u+u^{2^{s}}=1$, $v+v^{2^{2s}}=1,$ and $\beta\in\mathbb{F}_{2^{s}}$ satisfying $\beta=(t+1)^{-1}(\alpha^{2}+\alpha)+(t+1)(u^{2}+u)+1$, for some $\alpha\in\mathbb{F}_{2^{s}}$.
@@
$\begin{cases}\operatorname{DDT}_{F_{3}}(a,b)&\mbox{if }c=d=0,\\ \operatorname{DDT}_{F_{3}}(c,d)&\mbox{if}~{}ac\neq 0,a=c\mbox{ and }b=d;\mbox{or }a=0,b=0,c\neq 0;\mbox{ or }\\ &~{}\operatorname{DDT}_{F_{3}}(c,d)=4,a=z_{1}+z_{2}\mbox{ and }b=F_{3}(z_{1})+F_{3}(z_{2});\mbox{ or }\\ &~{}\operatorname{DDT}_{F_{3}}(c,d)=4,a=z_{1}+z_{2}+c\mbox{ and }b=F_{3}(z_{2})+F_{3}(z_{2}+c),\\ 0&\mbox{otherwise},\end{cases}$
@@
Case 2. Let $\frac{z_{2}}{c}=\frac{z_{1}}{c}+(t+1)u+\alpha+1$, then by the same argument as in Case 1, we have $a=c((t+1)u+\alpha+1)$ and $b=d(\alpha+(t+1)u^{2^{s}}+1)+F_{3}(c)((t+1)^{-1}(\alpha^{4}+\alpha^{2})+(t+1)^{3}(u^{4}+u^{2})+\alpha^{2}t+\alpha t+t(t+1)(u+u^{2})).$
@@
Next, we compute the UBCT and LBCT entries for the Bracken-Leander function. Using Corollary 7.1 the LBCT entries are straightforward to determine when $abc=0$, except the case where $a\neq b,ab\neq 0$ and $c=0$. Furthermore, if $ab\neq 0,a\neq b$ and $c=0$, the conditions derived for LBCT entries are identical to those obtained for the case where $abc\neq 0$ and $a\neq b$. Similarly, in the case of UBCT entries, we only need to focus on the case when $a\neq 0$ and $bc=0$ as the other cases are trivial if $abc=0$. The conditions in this case are identical to those for $abc\neq 0$. Therefore, it is sufficient to determine the UBCT and LBCT entries when $a,b,c,d\in\mathbb{F}_{2^{n}}^{*}$.
We then use (also its notations) of Corollary 7.2 to compute the LBCT and UBCT entries for the Bracken-Leander function when $abcd\neq 0$. As we have computed $\operatorname{EBCT}_{F_{3}}(a,b,c,d)=2$, when $a=c,b=d$ and $\operatorname{DDT}_{F_{3}}(c,d)=2$, then $\operatorname{LBCT}_{F_{3}}(a,c,d)=2$, for $a=c$ when $\operatorname{DDT}_{F_{3}}(c,d)=2$ and, $\operatorname{UBCT}_{F_{3}}(c,d,b)=2$, for $b=d$ when $\operatorname{DDT}_{F_{3}}(c,d)=2$.
@@
Via our Corollary 7.1, we can also give a very short proof for the EBCT, LBCT and UBCT entries of the inverse function for $n$ even, which is the main result in [13, 17] (for $n$ odd, see the previous section). Since $F_{4}$ is a permutation, the case for $abcd=0$ follows directly from Lemma 2.8. Therefore, we focus our discussion in the following corollary on the cases where $a,b,c,d\in\mathbb{F}_{2^{n}}^{*}$.
@@
Let $F_{4}(X)=X^{2^{n}-2}$ on $\mathbb{F}_{2^{n}}$, where $n$ is even. For $a,b,c,d\in\mathbb{F}_{2^{n}}^{*}$, we have
$\operatorname{EBCT}_{F_{4}}(a,b,c,d)$ =$=\begin{cases}4&\mbox{if }~{}a=c=\frac{1}{b}=\frac{1}{d};\mbox{ or }\\ &~{}b=\frac{1}{a},d=\frac{1}{c},\mbox{ and }(ad)^{2}+ad+1=0,\\ 2&\mbox{if}~{}a=c,b=d,d\neq\frac{1}{c}\mbox{ and }{\rm Tr}\left(\frac{1}{cd}\right)=0,\\ 0&\mbox{otherwise,}\end{cases}$
$\operatorname{LBCT}_{F_{4}}(a,b,c)$ =$=\begin{cases}4&\mbox{if }~{}a=b=\frac{1}{c};\mbox{ or }\\ &~{}b=\frac{1}{c},(ac)^{2}+ac+1=0,\\ 2&\mbox{if}~{}a=b\neq\frac{1}{c},{\rm Tr}\left(\frac{1}{bc}\right)=0,\\ 0&\mbox{otherwise,}\end{cases}$
$\operatorname{UBCT}_{F_{4}}(a,b,c)$ =$=\begin{cases}4&\mbox{if }~{}b=c=\frac{1}{a};\mbox{ or }\\ &~{}b=\frac{1}{a},(ac)^{2}+ac+1=0,\\ 2&\mbox{if}~{}c=b\neq\frac{1}{a},{\rm Tr}\left(\frac{1}{ab}\right)=0,\\ 0&\mbox{otherwise.}{}\end{cases}$
@@
If $a=c,b=d$, $d\neq\frac{1}{c}$ and ${\rm Tr}\left(\frac{1}{cd}\right)=0$, then $\operatorname{EBCT}_{F_{4}}(a,b,c,d)=\operatorname{DDT}_{F_{4}}(c,d)=2$. If $d=\frac{1}{c}$, and without loss of generality, we let $z_{1}=0,z_{2}=\omega c$. If $a=c,b=d$, then $\operatorname{EBCT}_{F_{4}}(a,b,c,d)=\operatorname{DDT}_{F_{4}}(c,d)=4$. If $a=z_{1}+z_{2}=\omega c$ and $b=F_{4}(z_{1})+F_{4}(z_{2})=\frac{1}{\omega c}$, then $\operatorname{EBCT}_{F_{4}}(a,b,c,d)=\operatorname{DDT}_{F_{4}}(c,d)=4$. If $a=z_{1}+z_{2}+c=\omega^{2}c$ and $b=F_{4}(z_{1})+F_{4}(z_{2}+c)=\frac{1}{\omega^{2}c}$, then $\operatorname{EBCT}_{F_{4}}(a,b,c,d)=\operatorname{DDT}_{F_{4}}(c,d)=4$. Note that these last two cases can be summarized as $b=\frac{1}{a}$, $d=\frac{1}{c}$ and $(ad)^{2}+ad+1=0$. In all other cases, $\operatorname{EBCT}_{F_{4}}(a,b,c,d)=0$.
@@
Clearly, when $d=\frac{1}{c}$, we can write $A=\{(c,c,d),(\omega c,c,d),(\omega^{2}c,c,d)\bigm{|}\operatorname{DDT}_{F_{4}}(c,d)=4\}$, as, in the case $\operatorname{DDT}_{F_{4}}(c,d)=4$, $(c,c,d)\in A$, for $b=d$, $(\omega c,c,d)\in A$, for $b=\frac{1}{\omega c}$ and $(\omega^{2}c,c,d)\in A$, for $b=\frac{1}{\omega^{2}c}$. Thus, from Corollary 7.2, we have $\operatorname{LBCT}_{F_{4}}(a,c,d)=4$, for $a\in\{c,\omega c,\omega^{2}c\}$ and $d=\frac{1}{c}$. This is equivalent to $\operatorname{LBCT}_{F_{4}}(a,c,d)=4$ when $d=\frac{1}{c}$ and $(ad)^{2}+ad=1=0$.
In a similar way, one can write $B=\{(c,d,d),(c,d,\frac{1}{\omega c}),(c,d,b=\frac{1}{\omega c})\bigm{|}\operatorname{DDT}_{F_{4}}(c,d)=4\}$, as, in the case $\operatorname{DDT}_{F_{4}}(c,d)=4$, $(c,d,d)\in B$, for $a=c$, $(c,d,\frac{1}{\omega c})\in B$, for $a=\omega c$ and $(c,d,\frac{1}{\omega c})\in B$, for $a=\omega^{2}c$, and the claim follows. ∎